# A Longitudinal, Multinational, and Multilingual Corpus of News Coverage of the Russo-Ukrainian War

### v1

@@ Abstract
We introduce dnipro, a novel longitudinal corpus of 246K news articles documenting the Russo-Ukrainian war from Feb 2022 to Aug 2024, spanning eleven media outlets across five nation states (Russia, Ukraine, U.S., U.K., and China) and three languages (English, Russian, and Mandarin Chinese). This multilingual resource features consistent and comprehensive metadata, and multiple types of annotation with rigorous human evaluations for downstream tasks relevant to systematic transnational analyses of contentious wartime discourse. dnipro’s distinctive value lies in its inclusion of competing geopolitical perspectives, making it uniquely suited for studying narrative divergence, media framing, and information warfare. To demonstrate its utility, we include use case experiments using stance detection, sentiment analysis, topical framing, and contradiction analysis of major conflict events within the larger war. Our explorations reveal how outlets construct competing realities, with coverage exhibiting polarized interpretations that reflect geopolitical interests. Beyond supporting computational journalism research, dnipro provides a foundational resource for understanding how conflicting narratives emerge and evolve across global information ecosystems.
@@ A Longitudinal, Multinational, and Multilingual Corpus of News Coverage of the Russo-Ukrainian War
Dikshya Mohanty, Taisiia Sabadyn, Jelwin Rodrigues, Chenlu Wang, Abhishek Kalugade, Ritwik Banerjee Stony Brook University Correspondence: {dimohanty, jelrodrigues, rbanerjee}@cs.stonybrook.edu
1 Introduction
@@
The Russo-Ukrainian war, which escalated on February 24, 2022 with the declaration of a "special military operation" by the Russian government, [^1] Within 48 hours, media in Russia were banned from describing the developments as “war” AFP (2022). has significantly impacted the global geopolitical landscape Derviş (2023); Silvertsen (2024). The war has redrawn traditional alliances, tested the resolve of international institutions, and created new geopolitical fault lines. Despite initial expectations of a swift Russian victory, Ukraine’s resistance has resulted in a protracted war with substantial loss of life and human capital Watling et al. (2023); Pancevski (2024). This conflict has also been characterized by a quieter battle for narrative control, with news coverage from several nations playing a pivotal role in shaping global public opinion and influencing the war’s trajectory Bradshaw et al. (2024). These sources reflect their national priorities, strategic perspectives, and the perceptions they intend to create among information consumers Johais and Meis (2024). Yet, despite the centrality of media and narrative framing, there remains a lack of comprehensive datasets for longitudinal analysis of how nations have framed the war for their domestic and international audiences.
@@
We introduce a collection of Diverse Narratives and International Perspectives on the Russo-Ukrainian Offensive (dnipro), a comprehensive longitudinal corpus of news coverage from February 2022 to August 2024222An expanded version is being developed, which will include news coverage up to February 2025 and provide a longitudinal 3-year corpus of this protracted war.—the longest time span of any multinational or multilingual news corpus on the Russia-Ukraine war. The corpus encompasses eleven influential media outlets and news agencies aligned with five primary geopolitical actors with strategic interests in the conflict. Following methodological approaches established by Nygren et al. (2016) and Roman et al. (2017), the sources include media from direct participants—Russia and Ukraine—and influential global powers with economic, diplomatic, or military involvement: USA, UK, and China. The inclusion of state-affiliated media from Russia and Ukraine capture opposing conflict narratives, while those from China, the USA, and the UK represent the narratives projected by nations with significant stakes in the conflict’s outcome, and significant ability to influence that outcome Zhang et al. (2016); Thussu et al. (2018).
@@
dnipro offers rich annotations for entity-level stance, sentiment, and topical framing, enabling event tracking and other longitudinal analyses across geopolitical perspectives. The dataset does not attempt to verify the accuracy of any claims; rather, it provides a resource for studying how wartime narratives are constructed and disseminated by major geopolitical actors. [^3] This aligns with established methodologies in discourse analysis Entman (1993); Fairclough (2013) examining how language and framing construct social reality. Further, Gamson and Modigliani (1989) showed how this view offers a powerful tool to understand the competition that shapes public understanding of contentious events. The corpus is uniquely suited to such studies, as it comprises 246,229 articles across 11 media sources and 3 languages, spanning 5 diverse nations, over a period of 31 months. Our key contributions include:
@@
    a longitudinal multinational and multilingual corpus of news coverage spanning the first 30 months of the Russia-Ukraine war;
@@
    annotations supporting comparative analyses of named entities, stance, and framing; and
@@
    the inclusion of diverse media sources enabling cross-cultural examination of geopolitical narratives.
Next, we place dnipro within the landscape of existing conflict-related corpora and highlight its distinct contributions to the field (§ 2), and describe how our design and construction choices (§ 3) make it a powerful resource for large-scale analyses of wartime narratives. We employ dnipro in several applications, using state-of-the-art language understanding systems as well as additional annotations (§ 4 and § 5), demonstrating its utility.
2 Related Work
@@
Twitter & Reddit-based datasets are among the largest. Chen and Ferrara (2023) aggregate over 570 million tweets in English, Russian, and Ukrainian to study information propagation and influence operations, while Zhu et al. (2022) provide 8 million posts correlating post categories with pro-Ukrainian stance. Shevtsov et al. (2022) introduce 57.3 million English tweets annotated for sentiment and hate speech, and the Invasion@Ukraine corpus provides 8.7 million English tweets for content moderation and conspiracy narrative analysis Susanne Pohl et al. (2023). The VoynaSlov collection Park et al. (2022) bears similarities to our work through sentence-level media framing labels, though they focus solely on Russian media.
Alyukov et al. (2023) introduce one of the first cross-platform social media datasets, integrating 1.7 million posts to enable propaganda and public opinion analyses. Telegram-specific corpora include 4.1 million posts with pro/anti-Kremlin channel tags Bawa et al. (2025) and 2.3 million posts with binary tags for human rights violations Nemkova et al. (2023).
@@
Hakimov and Cheema (2024) introduce a corpus of 1.5 million tweets in 60 languages with automated sentiment, stance, and named entity annotations, offering the widest multilingual social media dataset. The RUWA corpus Khairova et al. (2024) provides 16,500 news articles using semantic similarity for misinformation detection, but excludes the overarching war theme and curates only specific event-based articles. dnipro supersedes it by orders of magnitude in coverage, topics, multilinguality, temporality, and size. Ibrahim et al. (2025) offer longitudinal media analysis of two years of news data, but from nations with auxiliary involvement.
@@
Annotation practices differ widely: manual schemata for framing Park et al. (2022), channel stance Bawa et al. (2025), and human rights violations Nemkova et al. (2023); automated annotations for large-scale labeling Susanne Pohl et al. (2023); Hakimov and Cheema (2024). Most resources rely on social media identifiers that degrade quickly, impeding reproduction. Few provide longitudinally contiguous, multilingual, news-based content annotated for multiple tasks.
dnipro extends existing corpora by focusing on transnational and multilingual news coverage – underrepresented in prior social media-centered work – enabling comparative analyses of state-aligned narratives across geopolitical actors. It reflects news coverage as disseminated, without alteration, serving as a neutral resource for empirical studies. Our design choices spanning source selection, temporal coverage, and structured annotations make dnipro a powerful resource for large-scale automated analysis of wartime narratives (§ 5).
3 The dnipro Corpus
@@
Our dataset is constructed through a systematic multi-stage pipeline targeting news outlets and state-affiliated agencies selected on the basis of three criteria: (1) global influence and prominence within national media ecosystems, (2) sustained coverage of the Russo-Ukrainian war, and (3) discernible editorial positioning ranging from independent journalism to state-aligned messaging. To the extent possible, access to full article archives was obtained through public RSS feeds, institutional subscriptions, or source-specific data collection pipelines. However, not all media outlets provided consistent or unrestricted access to their archives, and some selections reflect practical constraints in addition to topical relevance.
3.1 Data Acquisition
dnipro is a corpus spanning 31 months of daily data collection, from February 1, 2022 to August 31, 2024. To identify articles with substantive coverage of the Russo-Ukrainian conflict, we employ a tiered strategy. For outlets with dedicated sections or labels related to the war, we collect all articles published under those categories. Where available, we also extract articles tagged with conflict-specific metadata (e.g., “Ukraine conflict”, “special military operation”, “Russia-Ukraine war”). For sources lacking such taxonomies, we apply keyword-based filtering to isolate relevant content. However, some sources impose significant restrictions on retrospective API-based keyword searches. In these cases, we obtain the older articles directly through web browsers, operating within the scope of respective terms of service and copyright restrictions. To ensure quality control, we manually validate stratified random samples across sources, languages, and time periods throughout the acquisition process.
Total no. of articles 246,229
@@ 3.1 Data Acquisition
No. articles/day (avg.) 261.1
No. of sources 11
Languages covered N/A
Temporal span Feb 1, 2022 - Aug 31, 2024
No. of non-English articles 61,637 (25.03%)
Geopolitical entities Russia, Ukraine, China, USA, UK
@@
Table 1: Key descriptive statistics of dnipro.
English-language “Western” media from the US and the United Kingdom, namely The Financial Times (FT) and Cable News Network (CNN), were included as part of the Western bloc, reflecting the perspectives of countries with significant political, economic, and military involvement in the conflict. These outlets serve domestic and international audiences and are central to understanding how Western-aligned narratives are constructed and disseminated. No additional filtering of articles is done by search queries as these media houses offer dedicated sections for content related to the Russo-Ukrainian war. We collect a total of 5,760 and 4,218 articles from CNN and FT, respectively.
@@
The Ukrainian perspective is represented by articles from Censor.net, The Kyiv Independent (KI), and European Pravda (EP), all of which provide English-language versions. While KI and EP shape Ukraine’s global image, Censor influences the domestic audience. Their reporting styles vary—from Censor’s rapid updates to EP’s in-depth opinions and interviews. As sources targeting the global Anglosphere, KI and EP offer sections dedicated to the Russo-Ukrainian War: “War” and “War with Russia”, respectively. Censor’s news reporting approach largely covers the war, evidenced by the majority fraction of Ukrainian war-related tags in their “News” section. As such, no keyword search queries are applied to any of the Ukrainian sources. The collection from these courses amount to 19,706 articles from Kyiv Independent, 2,212 from European Pravda, and 31,748 from Censor.
@@
article_id A unique UUID v4 identifier for the record.
url The persistent URL from which the article was sourced.
publication_date The article’s publication date in ISO 8601 format (yyyy-mm-dd).
source_name The normalized name of the news outlet (e.g., Xinhua).
source_country The country of origin for the news outlet, in ISO 3166-1 alpha-3.
language The article language, in ISO 639-1.
title The article’s title (i.e., headline).
article_length The number of words in the cleaned article text (post processing).
collection_date The date (in ISO 8601 format) when the URL was accessed.
Table 2: Core metadata fields for each article.
3.2 Metadata Schema and Statistics
The core of dnipro is a collection of rich, consistent metadata records for 246,229 news articles. Its key descriptive statistics are shown in Table 1. Each record is described by a unified schema (described in Table 2) designed to facilitate large-scale longitudinal and comparative analyses.
The overall composition of the natural language data—both in terms of the number of news articles and their sizes—across individual news sources and aggregated by geopolitical entities, is presented in Figure 1.
3.3 Data Records
dnipro is publicly available on Zenodo under a CC-BY-NC 4.0 license. [^4] The data, along with all processing code and annotation prompts, will be released upon acceptance. It is structured for ease of use and distributed as a single, compressed tar.bz2 archive containing a primary metadata file, annotation files, and a documentation file. [^5] We do not directly include the texts due to copyright restrictions. The provided URLs serve as persistent identifiers enabling reproduction of dnipro, like most social media datasets. If an article becomes unavailable, the collection_date and url enable retrieval via the Internet Archive’s WayBack Machine.
@@ 3.3 Data Records
    metadata.parquet contains the complete set of 246,229 metadata records as described in § 3.2. The Parquet format was chosen for its efficient compression and columnar storage, which enables fast querying and analysis using popular data science tools.
@@
    sentiments.parquet provides sentiment scores for select entities and events across multiple news sources. For politically important and frequently occurring entities (e.g., Ukraine, Russia, Zelensky, Kherson), these scores are aggregated monthly. For entities tied to key events within this war (e.g., Bucha, Kursk, Zaporizhzhia), they are aggregated daily.
@@
    topical-frames.parquet provides one row per article_id with its corresponding topical frames anchored in the nine-category schema described in Table 3.

### v2

@@ Abstract
We present dnipro, a corpus of 246K news articles from the Russo-Ukrainian war (Feb 2022 – Aug 2024) spanning eleven outlets across five nation-states (Russia, Ukraine, U.S., U.K., China) and three languages. The corpus features comprehensive metadata and human-evaluated annotations for stance, sentiment, and topical framing, enabling systematic analysis of competing geopolitical narratives. It is uniquely suited for empirical studies of narrative divergence, media framing, and information warfare. Our exploratory analyses reveal how media outlets construct incompatible realities through divergent attribution and topical selection without direct refutation of opposing narratives. dnipro empowers empirical research on narrative evolution, cross-lingual information flow, and computational detection of implicit contradictions in fragmented information ecosystems.
Keywords: multilingual corpus, media framing, geopolitical narratives, longitudinal analysis, information warfare
[…]
@@ A Longitudinal, Multinational, and Multilingual Corpus of News Coverage of the Russo-Ukrainian War
Dikshya Mohanty1, Taisiia Sabadyn1, Jelwin Rodrigues1,
Chenlu Wang1, Abhishek Kalugade1, Ritwik Banerjee1,2
1Department of Computer Science, 2AI Innovation Institute
Stony Brook University, New York
{dimohanty, jelrodrigues, chenlwang, rbanerjee}@cs.stonybrook.edu
{taisiia.sabadyn, abhishek.kalugade}@stonybrook.edu
Abstract content
1.  Introduction
@@
The Russo-Ukrainian war, which escalated on February 24, 2022 with the declaration of a “special military operation” by the Russian government, [^1] Within 48 hours, media in Russia were banned from describing the developments as “war” AFP (2022). has significantly impacted the global geopolitical landscape Derviş (2023); Silvertsen (2024). The war has redrawn traditional alliances, tested the resolve of international institutions, and created new geopolitical fault lines. Despite initial expectations of a swift Russian victory, Ukraine’s resistance has resulted in a protracted war with substantial loss of life and human capital Watling et al. (2023); Pancevski (2024). This conflict has also been characterized by a quieter battle for narrative control, with news coverage from several nations playing a pivotal role in shaping global public opinion and influencing the war’s trajectory Bradshaw et al. (2024). These sources reflect their national priorities, strategic perspectives, and the perceptions they intend to create among information consumers Johais and Meis (2024). Yet, despite the centrality of media and narrative framing, there remains a lack of comprehensive datasets for longitudinal analysis of how nations have framed the war for their domestic and international audiences.
@@
We introduce a collection of Diverse Narratives and International Perspectives on the Russo-Ukrainian Offensive (dnipro), a comprehensive longitudinal corpus of news coverage from February 1, 2022 to August 31, 2024 — the longest time span of any multinational or multilingual news corpus on the Russia-Ukraine war. The code-base and supporting materials are publicly available. The corpus encompasses eleven influential media outlets and news agencies aligned with five primary geopolitical actors with strategic interests in the conflict. Following methodological approaches established by Nygren et al. (2016) and Roman et al. (2017), the sources include media from direct participants—Russia and Ukraine—and influential global powers with economic, diplomatic, or military involvement: the USA, the UK, and China. The inclusion of state-affiliated media from Russia and Ukraine captures opposing conflict narratives that play a pivotal role in shaping assessments under uncertain information Saletta et al. (2020). Media from China, the USA, and the UK represent perspectives of nations with substantial geopolitical and economic interests in the conflict’s trajectory, as well as the military, diplomatic, and economic leverage to influence its outcome Zhang et al. (2016); Thussu et al. (2017). These diverse perspectives and sources enable more comprehensive analysis of complex developments by helping analysts overcome the limitations of single-narrative framings Werd (2018).
@@
dnipro offers rich annotations for topical framing, entity-level stance, and entity-level sentiment, enabling event tracking and other longitudinal analyses across geopolitical perspectives. The dataset does not attempt to verify the accuracy of any claims; rather, it provides a resource to study how wartime narratives are constructed and disseminated by major geopolitical actors. [^2] This aligns with established methodologies in discourse analysis Entman (1993); Fairclough (2010) examining how language and framing construct social reality. Gamson and Modigliani (1989) demonstrated how this approach reveals the competition shaping public understanding of contentious events. The corpus is uniquely suited to such studies, as it comprises 246,229 articles across 11 media sources and 3 languages, spanning 5 diverse nations, over a period of 31 months. Our key contributions include:
@@
    a longitudinal, multinational, and multilingual corpus of news coverage spanning the first 31 months of the Russia-Ukraine war;
@@
    inclusion of diverse media, enabling cross-cultural examination of conflict narratives; and
@@
    annotations for comparative analyses of named entities, sentiment, stance, and topical framing.
Next, we place dnipro within the landscape of existing conflict-related corpora to highlight its distinct contributions to the field (§ 2), and describe how our design and construction choices (§ 3) make it a powerful resource for large-scale analyses of wartime narratives. We employ dnipro in several applications, using state-of-the-art natural language understanding systems and additional annotations (§ 4 and § 5), demonstrating its utility.
2.  Related Work
@@
Datasets based on Twitter/X and Reddit are among the largest. Chen and Ferrara (2023) aggregate over 570 million tweets in English, Russian, and Ukrainian to study information propagation and influence operations, while Zhu et al. (2022) provide 8 million posts correlating post categories with pro-Ukrainian stance. Shevtsov et al. (2022) introduce 57.3 million English tweets annotated for sentiment and hate speech, and the Invasion@Ukraine corpus provides 8.7 million English tweets for content moderation and conspiracy narrative analysis Susanne Pohl et al. (2023). The VoynaSlov collection Park et al. (2022) bears similarities to our work through sentence-level media framing labels, though it focuses solely on Russian media. Alyukov et al. (2023) introduce one of the first cross-platform social media datasets, integrating 1.7 million posts to enable the analysis of propaganda and public opinion. Telegram-specific corpora include 4.1 million posts with pro/anti-Kremlin channel tags Bawa et al. (2025) and 2.3 million posts with binary tags for human rights violations Nemkova et al. (2023). Hakimov and Cheema (2024) introduce a corpus of 1.5 million tweets in 60 languages with automated annotations for sentiment, stance, and named entities. Their corpus is one of the widest multilingual social media datasets currently available.
@@
News-based corpora are comparatively sparse. The RUWA corpus Khairova et al. (2024) provides 16,500 news articles using semantic similarity for misinformation detection, but excludes the overarching war theme and curates only specific event-based articles. dnipro supersedes it by orders of magnitude in coverage, topics, multilinguality, temporality, and size. A longitudinal media analysis corpus is provided by Ibrahim et al. (2025), with two years of news data, but its coverage is restricted to nations with auxiliary involvement in the Russo-Ukrainian war.
@@
Annotation practices across existing resources differ widely: manual schemata for framing Park et al. (2022), channel stance Bawa et al. (2025), and human rights violations Nemkova et al. (2023); as well as automated large-scale labeling Susanne Pohl et al. (2023); Hakimov and Cheema (2024). Most resources rely on social media identifiers that degrade quickly, impeding reproduction.
dnipro addresses these gaps by providing longitudinally contiguous, multilingual news coverage from five primary geopolitical actors (Russia, Ukraine, U.S., U.K., and China), with comprehensive annotations for multiple downstream tasks. Unlike social media-focused datasets optimized for volume and engagement patterns, dnipro enables systematic comparative analyses of state-aligned narrative framing across languages, nations, and outlets, over a duration of 2 years and 7 months, making it a unique and powerful resource for empirical studies on the creation and dissemination of competing wartime realities (§ 5).
3.  The dnipro Corpus
We present dnipro, a longitudinal corpus of 246,229 news articles spanning 31 months of the Russo-Ukrainian war (February 1, 2022 – August 31, 2024). Our corpus captures competing narratives from 11 outlets across five nation states (Russia, Ukraine, U.S., U.K., and China) and three languages (English, Russian, and Mandarin Chinese). Its key descriptive statistics appear in Table 1.
@@ 3.  The dnipro Corpus
Each article includes structured metadata following a unified schema (shown in Table 2) designed to facilitate large-scale longitudinal and comparative analyses across sources, languages, and geopolitical contexts. dnipro is constructed through a systematic multi-stage pipeline targeting news outlets and state-affiliated agencies selected on the basis of three criteria: (i) global influence and prominence within national media ecosystems, (ii) sustained coverage of the Russo-Ukrainian war, and (iii) discernible editorial positioning ranging from independent journalism to state-aligned messaging. To the extent possible, access to full article archives was obtained through public RSS feeds, institutional subscriptions, or source-specific data collection pipelines. However, not all media outlets provided consistent or unrestricted access to their archives, and some selections reflect practical constraints in addition to topical relevance.
We first detail our collection methodology (§3.1), then describe the structure and access information for the data records (§3.2).
Corpus Scale
Total articles 246,229
Avg. no. of articles/day 261.1
@@
Coverage
Temporal span Feb 1, 2022 – Aug 31, 2024
No. of news sources 11 sources
Geopolitical entities Russia, Ukraine, USA, UK, China
Linguistic Composition
Languages covered English, Russian, Mandarin Chinese
Non-English articles 61,637 (25.03%)
@@
Table 1: dnipro at a glance: scale, temporal and geographic coverage, and multilingual composition.
3.1.  Data Acquisition
dnipro comprises 31 months of continuous daily data collection (Feb. 1, 2022 – Aug. 31, 2024). To identify articles with substantive coverage of the Russo-Ukrainian conflict, we employ a tiered strategy. For outlets with dedicated sections or labels related to the war, we collect all articles published under those categories. Where available, we also extract articles tagged with conflict-specific metadata (e.g., “Ukraine conflict”, “special military operation”, “Russia-Ukraine war”). For sources lacking such taxonomies, we apply keyword-based filtering to isolate relevant content. However, some sources impose significant restrictions on retrospective API-based keyword searches. In these cases, we obtain the older articles directly through web browsers, operating within the scope of respective terms of service and copyright restrictions. To ensure quality control, we manually validate stratified random samples across sources, languages, and time periods throughout the acquisition process.
Field Description
Identifiers & Source
article_id Unique UUID v4 identifier
url URL of the article
source_name Normalized source name (e.g., Xinhua)
source_country Country of origin (ISO 3166-1 alpha-3)
Content & Language
title Article headline
article_length Word count (post-processing)
language Article language (ISO 639-1)
Temporal Information
publication_date Publication date (ISO 8601)
collection_date Date of URL access (ISO 8601)
Table 2: Core metadata schema for dnipro articles.
English-language “Western” media from the United States and the United Kingdom, namely The Financial Times (FT) and Cable News Network (CNN), were included as part of the Western bloc, reflecting the perspectives of countries with significant political, economic, and military involvement in the conflict. These outlets serve domestic and international audiences, and offer central platforms to help us understand how Western-aligned narratives are constructed and disseminated. No additional filtering of articles is done by search queries as these media houses offer dedicated sections for content related to the Russo-Ukrainian war. We collect a total of 5,760 and 4,218 articles from CNN and FT, respectively.
@@ 3.1.  Data Acquisition
The Ukrainian perspective is represented by articles from Censor.net, The Kyiv Independent, and European Pravda, all of which provide English-language versions. While Censor.net influences Ukraine’s domestic audience, the other two shape her global image. Their reporting styles vary—from Censor’s rapid updates to the in-depth opinions and interviews found in European Pravda. As sources targeting the global Anglosphere, The Kyiv Independent and European Pravda offer sections dedicated to the Russo-Ukrainian War, labeled respectively with “War” and “War with Russia”. The news reporting approach of Censor.net largely covers the war, evidenced by the majority fraction of Ukrainian war-related tags in their “News” section. As such, no keyword search queries are applied to any of the Ukrainian sources. The collection from these sources amounts to 19,706 articles from The Kyiv Independent, 2,212 from European Pravda, and 31,748 from Censor.net.
@@
Figure 1 shows the distribution of articles and their average lengths across individual news sources and aggregated by geopolitical entities, providing an overview of the varying scales and styles of coverage across geopolitical perspectives.
3.2.  Data Records
The dnipro corpus is available at no cost for non-commercial use under the CC-BY-NC 4.0 license, [^3] The complete corpus is available on Zenodo: doi.org/10.5281/zenodo.18470677 and the corresponding processing scripts and annotation prompts are available on GitHub.444github.com/dikshyam/dnipro-codebase It is structured for ease of use and distributed as separate downloadable files, comprising a primary metadata file, annotation files, and a documentation file. Due to copyright restrictions, we cannot directly include the texts. The provided URLs, however, serve as persistent identifiers enabling reproduction of dnipro, like most social media datasets. If an article becomes unavailable, the collection_date and url fields enable retrieval via the Internet Archive’s WayBack Machine.555web.archive.org
@@ 3.2.  Data Records
    metadata.parquet contains the complete set of 246,229 metadata records as described in Table 2. The Parquet format was chosen for its efficient compression and columnar storage, which enables fast querying and analysis using popular data science tools.
@@
    topical-frames.parquet provides one row per article_id with its corresponding topical frames anchored in categories proposed by Habermas (1991) (Table 3).
@@
    sentiments.parquet provides sentiment scores for select entities and events across multiple news sources. For politically important and frequently occurring entities (e.g., “Ukraine”, “Russia”, “Zelensky”, “Kherson”), these scores are aggregated monthly. For entities tied to key events within this war (e.g., “Bucha” or “Zaporizhzhia”), they are aggregated daily. The events used in this work are described in § 5 (Table 4).